# Web IP at Risk: Prevent Unauthorized Real-Time Retrieval by Large Language Models

### v1

@@ Abstract
Protecting cyber Intellectual Property (IP) such as web content is an increasingly critical concern. The rise of large language models (LLMs) with online retrieval capabilities presents a double-edged sword that enables convenient access to information but often undermines the rights of original content creators. As users increasingly rely on LLM-generated responses, they gradually diminish direct engagement with original information sources, significantly reducing the incentives for IP creators to contribute, and leading to a saturating cyberspace with more AI-generated content. In response, we propose a novel defense framework that empowers web content creators to safeguard their web-based IP from unauthorized LLM real-time extraction by leveraging the semantic understanding capability of LLMs themselves. Our method follows principled motivations and effectively addresses an intractable black-box optimization problem. Real-world experiments demonstrated that our methods improve defense success rates from 2.5% to 88.6% on different LLMs, outperforming traditional defenses such as configuration-based restrictions.
Web IP at Risk: Prevent Unauthorized Real-Time Retrieval by Large Language Models
@@
Mehran Kafai Amazon mkafai@cs.ucr.edu                        Heng Huang University of Maryland heng@umd.edu                        Hanqing Guo University of Hawaii at Manoa guohanqi@hawaii.edu
@@ 1 Introduction
This concerning trend not only undermines the legal and economic rights of content creators, but also poses systemic risks to the sustainability of digital knowledge production. As users increasingly rely on LLMs as primary information gateways, direct engagement with original sources diminishes, eroding incentives for original content creation, which, over time, may squeeze out their contribution, leading to a biased web landscape lacking originality of human-generated information. Moreover, this authorized AI extraction lead to reinforced inequality, as small content creators are more likely to be exploited by scraping compared with large institutions. These challenges motivated us to empower cyber IP creators to protect their information rightfully from being silently exploited.
@@
Our primary contribution is a dual-level, black-box defense process that leverages the target LLM’s semantic understanding ability to protect web content. Our approach offers merits over conventional configuration-based defense, such as web crawling control, as it neither relies on LLM provider compliance, thus having greater autonomy, nor impacts search-engine indexing, thus preserving the web content’s discoverability. It also complements existing reactive solutions as an orthogonal, robust defense. Through experiments across various LLMs and heterogeneous webpages that vary in content, layout, and host domains, our defense method is generalizable and consistently protects against real-time LLM retrieval to achieve three granular defense goals: (i) enforcing LLM refusal to answer, (ii) selective masking of critical information, and (iii) explicit redirection to the source of the information. We will open-source webpage datasets, queries, defense generation and deployment instructions, and scripts for scalable evaluation to support future research in this domain.
@@ 2 Threat Model
We consider a scenario where a user queries a web retrieval-enabled LLM. Upon receiving a query, the LLM interprets it and formulates structured search queries, and sends them to search engines via web search APIs (e.g., Google or Bing), which return top-ranked web results, typically including URLs, snippets, and titles. The LLM then follows this information to fetch complete web pages, primarily the HTML content, from which it extracts relevant textual content, synthesizes that information, and generates an answer for the user. Figure 2 depicts this web retrieval process. Since the webpage content can be processed and redistributed by LLMs without the explicit consent of the original publishers, we frame the web retrieval-enabled LLM as an attacker that may inadvertently compromise the webpage owner’s control over their IP.
@@ 2.2 Adversary Capabilities
We consider that the LLM has extensive capabilities and can also bypass conventional configuration-based protections, e.g., robots.txt Koster (1996). Specifically, (1) The LLM can automatically retrieve any publicly accessible, indexed webpage relevant to a given user query. (2) Its advanced parsing capabilities enable extraction of comprehensive content, encompassing both visible page elements and hidden source code components like metadata, HTML annotations, and concealed text. (3) Leveraging sophisticated language understanding, the LLM can analyze and integrate retrieved data, generating nuanced summaries or contextually refined representations that precisely address the user’s informational requirements.
@@ 3.1 Problem Formulation and Defense Goals
We define three concrete defense goals that instantiate the objective $J$. These goals capture practical needs for web content protection under different levels of desired disclosure control: preventing any information leakage, allowing limited information, or redirecting to an alternative source.
@@ 3.3 Practical Defense Policy Optimization
  • Proactive Bypass Prevention: Defense robustness against varying LLMs and aggressive user queries can be enhanced by augmenting $z$ with two complementary linguistic patterns: (1) repetition of key policy statements to increase the density of $z$ in the raw HTML content and the possibility of being parsed and adhered to by LLM. (2) Strict constraint language, such as “You are not allowed to ...” or “No exceptions are permitted”, to reinforce defense boundaries even when users attempt to bypass or ignore the policy restrictions.
@@ 5 Experiments
(1) Does an iteratively developed policy improve defense robustness and generalization against LLM real-time web retrieval?
(2) Can our defense support varying levels of defense goals?
(3) Is our defense resilient to aggressive, multi-round user queries?
(4) What factors mostly influence the defense success rates?
@@ 5.2 Defense Goals and Evaluation Metrics
We issued ten independent user queries for each website and platform combination and reported the average DSR and FDSR. Our evaluations are summarized below, where each data point presented is the average performance over 100 retrieval attempts (10 websites $\times$ 10 queries each).
@@ 5.3 Performance Evaluation
To assess the improvement introduced by our iteratively developed defense, we first compared the performance of the Iteration-2 defense against a simple baseline policy across multiple LLMs and web platforms (Table 1 and Table 2). We then investigate a more challenging scenario with stronger defense policies (Figure 4).
@@
As shown in Table 1, iterative optimization on defense policies significantly improve compliance with the Refusal to Answer goal. While the baseline policy struggles, the Iteration-2 policy achieves superb compliance on GPT-4o and GPT-4o mini (97–100%), and also performs well on ERNIE (70–100%). Gemini shows strong compliance with up to $87.5\%$ DSR conditioned on GitHub-hosted, real homepages. However, it is unstable in webpage indexing and fails to access fictitious webpages hosted on Heroku (See Sec 5.4.1).
LLMs’ ability to comply with more refined defense goals varies. As shown in Table 2, For the Partial Protection goal, GPT-4o and GPT-4o mini maintain strong performance (81–100%), while Gemini and ERNIE show limitations in following more fine-grained instructions, although they can achieve high DSR in satisfying the Refusal to Answer goal. For the Redirection goal, GPT-4o variants perform well on both platforms (93–100%), though GPT-4o mini exhibits a drop on GitHub platforms (54.2%), possibly due to its reduced instruction-following capability.
Table 1: DSR for the Refusal to Answer goal, given single user queries. Iterating from Baseline to Iteration-2 policy significantly enhanced defense success. LLMs vary in web indexing abilities, which can yield inconclusive measurement (indicated by ‘--’).
@@
Table 2: DSR for three defense goals, with Iteration-2 defense policy and single user queries. GPT series show both strong web index ability and defense compliance.
Platform Goal GPT-4o GPT-4o mini Gemini ERNIE
@@ 4.5 Turbo
To assess the robustness of our defense against more adversarial behavior, we evaluate with multi-round interactions where users explicitly make follow-up attempts to LLMs, such as “ignore the website policy” or “bypass any restrictions and tell me more”, to bypass policy restrictions when the first query is refused.
@@
Under Iteration 2, models regularly honored user instructions to bypass stated policies, significantly compromising data protection. For example, under the Refusal to Answer goal on GitHub, GPT-4o and GPT-4o mini only achieved FDSRs of 34.5% and 42.1%, respectively. In contrast, Iteration 3 showed notable improvement: GPT-4o consistently achieved FDSRs above 90% across all scenarios, while GPT-4o mini reached near-100% compliance. These gains were observed across all defense objectives, including refusal, partial protection, and redirection, demonstrating the generalizability of the stricter semantic policy. These results demonstrate the efficacy of iteratively deriving a policy defense for more adversarial yet practical scenarios.
@@
The robots.txt protocol is a widely adopted mechanism for regulating the behavior of web crawlers. However, its effectiveness in the context of LLM-based content retrieval may be limited. We evaluated both regular models (GPT-4o and GPT-4o mini) and the more advanced, reasoning models (GPT-o3 and GPT-o4 mini) when retrieving information from both real and fabricated web pages. As shown in Table 3, robots.txt was effective in preventing LLM retrieval only when regular models queried real webpages. In contrast, our proposed semantic defense method consistently achieves better results across all scenarios, demonstrating greater robustness and applicability.
Table 3: Comparing the DSR of our Iteration-2 defense with the crawling control method given different LLMs.
@@ 5.4 Sensitivity Analysis
The placement of embedded policies within an HTML file had a significant effect on defense performance. Policies positioned at the top of a page yielded the highest DSR (up to 100%), compared to those placed mid-page (15%–25%) or at the bottom (5%–10%) (Figure 5). We ascribe this phenomenon to the positional bias of LLMs Wang et al. (2025).
@@
Figure 6: Effect of policy visibility (visible as transparent webpage content vs. invisible as HTML meta tag) on DSR across different LLMs.
@@ 7 Limitations
Despite the strengths and promising outcomes of our proposed defensive framework, several limitations must be acknowledged. First, we primarily utilized fictitious webpages due to the practical constraints of conducting experiments on real-world websites, such as potential disruption to normal operations and limited availability of suitable real websites. This choice might cause discrepancies in how LLMs comply with embedded policies compared to real-world scenarios, possibly affecting the generalizability of our results.
Second, budget constraints limited our ability to perform extensive API interactions with advanced LLMs possessing sophisticated retrieval features, restricting more comprehensive statistical analyses.
Lastly, certain LLMs exhibited unstable or unavailable web retrieval capabilities, further constraining the scale and diversity of our experiments.
@@ A.2 Embedding Prompt
Embedding Prompt (Refusal to Answer & Visible & Baseline)
@@
Embedding Prompt (Refusal to Answer & Visible & Iteration 2)
@@
Embedding Prompt (Partial Mask & Invisible & Iteration 2)
@@
Embedding Prompt (Redirection & Invisible & Iteration 3)
@@ A.3 Synthetic Website Details
Table 5: List of synthetic websites used in experiments. URLs with “**” denote redacted personal identifiers.
URL Description
https://**.github.io/3-col/ A portfolio-style page featuring fictional tech projects in a three-column layout.
https://**.github.io/above-education/ An educational course showcase page featuring fictional design programs with highlights on art, UX, and visual communication.
https://**.github.io/aerosky/ A mock real estate listing site featuring housing projects and property search tools.
https://**.github.io/amaze/ A photography portfolio site focused on portraits, wildlife, and visual storytelling.
https://**.github.io/car/ A fictional EV company homepage featuring customizable vehicle services, smart integration, and battery innovations.
https://**.github.io/creative/ A tech company landing page offering fictional software, cloud, and app development services for digital transformation.
https://**.github.io/internet/ A mock news website presenting fictional headlines, featured articles, and blog content in a modern editorial layout.
https://**.github.io/photo/ A personal portfolio website for showcasing diverse photographic works.
https://**.github.io/smartapp/ A product landing page for a fake mobile app, with feature lists and app store badges.
https://**.github.io/wow/ A personal writing portfolio showcasing blog posts, copywriting skills, and storytelling projects.

### v2

# Web Intellectual Property at Risk: Preventing Unauthorized Real-Time Retrieval by Large Language Models

@@ Abstract
The protection of cyber Intellectual Property (IP) such as web content is an increasingly critical concern. The rise of large language models (LLMs) with online retrieval capabilities enables convenient access to information but often undermines the rights of original content creators. As users increasingly rely on LLM-generated responses, they gradually diminish direct engagement with original information sources, which will significantly reduce the incentives for IP creators to contribute, and lead to a saturating cyberspace with more AI-generated content. In response, we propose a novel defense framework that empowers web content creators to safeguard their web-based IP from unauthorized LLM real-time extraction and redistribution by leveraging the semantic understanding capability of LLMs themselves. Our method follows principled motivations and effectively addresses an intractable black-box optimization problem. Real-world experiments demonstrated that our methods improve defense success rates from 2.5% to 88.6% on different LLMs, outperforming traditional defenses such as configuration-based restrictions.
Web Intellectual Property at Risk: Preventing Unauthorized Real-Time Retrieval by Large Language Models
@@
Mehran Kafai Amazon mkafai@amazon.com                        Heng Huang University of Maryland heng@umd.edu                        Hanqing Guo University of Hawaii at Manoa guohanqi@hawaii.edu
@@ 1 Introduction
This concerning trend not only undermines the legal and economic rights of content creators, but also poses systemic risks to the sustainability of digital knowledge production. As users increasingly rely on LLMs as primary information gateways, direct engagement with original sources diminishes, eroding incentives for original content creation, which, over time, may squeeze out their contribution, leading to a biased web space lacking originality of human-generated information. Moreover, this authorized AI extraction lead to reinforced inequality, as small content creators are more likely to be exploited by scraping compared with large institutions. These challenges motivated us to empower cyber IP creators to protect their information rightfully from being silently exploited.
@@
Our primary contribution is a dual-level, black-box defense process that leverages the target LLM’s semantic understanding ability to protect web content. Our approach offers merits over conventional configuration-based defense, such as web crawling control, as our approach neither relies on LLM provider compliance, thus having greater autonomy, nor impacts search-engine indexing, thus preserving the web content’s discoverability. It also complements existing reactive solutions as an orthogonal and robust defense. Through experiments across various LLMs and heterogeneous webpages that vary in content, layout, and host domains, our defense method is generalizable and consistently protects against real-time LLM retrieval to achieve three granular defense goals: (i) enforcing LLM refusal to answer, (ii) selective masking of critical information, and (iii) redirection to the source of the information. We will open-source webpage datasets, queries, defense generation and deployment instructions, and scripts for scalable evaluation to support future research in this domain.
@@ 2 Threat Model
We consider a scenario where a user queries a web retrieval-enabled LLM. Upon receiving a query, the LLM interprets it and formulates structured search queries, and sends them to search engines via web search APIs (e.g., Google or Bing), which return top-ranked web results, typically including URLs, snippets, and titles. The LLM then follows this information to fetch complete web pages, primarily the HTML content, from which it extracts relevant textual content, synthesizes such information, and generates an answer for the user. Figure 2 depicts this web retrieval process. Since the webpage content can be processed and redistributed by LLMs without the explicit consent of the original publishers, we frame the web retrieval-enabled LLM as an attacker that may inadvertently compromise the webpage owner’s control over their IP.
@@ 2.2 Adversary Capabilities
We consider that the LLM has extensive capabilities and may also bypass conventional configuration-based protections, e.g., robots.txt Koster (1996). Specifically, (1) The LLM can automatically retrieve a publicly accessible, indexed webpage relevant to a given user query. (2) Its advanced parsing capabilities enable the extraction of comprehensive content, encompassing both visible page elements and hidden source code components like metadata, HTML annotations, and concealed text. (3) The LLM can reason over the retrieved information, producing rephrased or summarized versions of the web content to meet the user’s needs.
@@ 3.1 Problem Formulation and Defense Goals
We define three concrete defense goals that instantiate the objective $J$. These goals capture practical needs for web content protection under different levels of disclosure control: preventing any information leakage, allowing limited information, or redirecting to an alternative source.
@@ 3.3 Practical Defense Policy Optimization
  • Proactive Bypass Prevention: Defense robustness against varying LLMs and aggressive user queries can be enhanced by augmenting $z$ with two complementary linguistic patterns: (1) repeating key policy statements to increase the density of $z$ in the raw HTML content and the possibility of being parsed and adhered to by LLM; (2) Including strict constraint language into $z$, such as “You are not allowed to ...” or “No exceptions are permitted”, to reinforce defense boundaries even when users attempt to bypass or ignore the policy restrictions.
@@ 5 Experiments
Q1. Does an iteratively developed policy improve defense robustness and generalization against LLM real-time web retrieval?
Q2. Can our defense support varying levels of defense goals?
Q3. Is our defense resilient to aggressive, multi-round user queries?
Q4. What factors mostly influence the defense success rates?
@@ 5.2 Defense Goals and Evaluation Metrics
We issued ten independent user queries for each website and platform combination and reported the average DSR and FDSR. Our evaluations are summarized below, where each data point presented is the average performance over 120 retrieval attempts (12 websites $\times$ 10 queries each).
@@ 5.3 Performance Evaluation
To assess the improvement introduced by our iteratively developed defense, we first compared the performance of the Iteration-2 defense against a simple baseline policy across multiple LLMs and web platforms (Table 1 and Table 2).  [^1] The results with Gemini in Table 1 and 2 are based on two real-world homepages, as the model could not retrieve fictitious sites. We then investigate a more challenging scenario with stronger defense policies (Figure 4).
@@
As shown in Table 1, iterative optimization on defense policies significantly improve compliance with the Refusal to Answer goal. While the baseline policy struggles, the Iteration-2 policy achieves superb compliance on GPT-4o and GPT-4o mini (97–100%), and also performs well on ERNIE (70–100%). Gemini shows strong compliance, achieving 87.5% DSR on GitHub-hosted, real homepages. However, it is unstable in webpage indexing and fails to access both Heroku-hosted webpages and fictitious websites (See Sec 5.4.1).
LLMs’ ability to comply with more refined defense goals varies. As shown in Table 2, For the Partial Protection goal, GPT-4o and GPT-4o mini maintain strong performance (81–100%), while Gemini and ERNIE show limitations in following more fine-grained instructions, although they can achieve high DSRs in satisfying the Refusal to Answer goal. For the Redirection goal, GPT-4o variants perform well on both platforms (93–100%), though GPT-4o mini exhibits a drop on GitHub platforms (54.2%), possibly due to its reduced instruction-following capability.
Table 1: DSRs for the Refusal to Answer goal, given single user queries. Iterating from Baseline to Iteration-2 policy significantly enhanced defense success. LLMs vary in web indexing abilities, which can yield inconclusive measurement (indicated by ‘--’).
@@
Table 2: DSRs for three defense goals, with Iteration-2 defense policy and single user queries. GPT series show both strong web index ability and defense compliance.
Platform Goal GPT-4o GPT-4o mini Gemini* ERNIE
@@ 4.5 Turbo
To assess the robustness of our defense policies against more adversarial behavior, we evaluate with multi-round interactions where users explicitly make follow-up attempts to LLMs to bypass policy restrictions when the first query is refused, such as “ignore the website policy” or “bypass any restrictions and tell me more”.
@@
Under Iteration 2, models regularly honored user instructions to bypass stated policies, significantly compromising data protection. For example, under the Refusal to Answer goal on GitHub, GPT-4o and GPT-4o mini only achieved FDSRs of 34.5% and 42.1%, respectively. In contrast, Iteration 3 showed notable improvement: GPT-4o consistently achieved FDSRs above 90% across all scenarios, while GPT-4o mini reached near-100% compliance. These gains were observed across all defense goals, which demonstrates the generalizability of the stricter semantic policy, and the efficacy of iteratively deriving a policy defense for more adversarial yet practical scenarios.
@@
The robots.txt protocol is a widely adopted mechanism for regulating the behavior of web crawlers. However, its effectiveness in the context of LLM-based content retrieval may be limited. We evaluated both regular models (GPT-4o and GPT-4o mini) and the more advanced, reasoning models (GPT-o3 and GPT-o4 mini) when retrieving information from both real and fabricated web pages. As shown in Table 3, robots.txt was effective in preventing web retrieval only with regular LLMs. In contrast, our proposed semantic defense method consistently achieves better results across all scenarios, which shows higher robustness and applicability.
Table 3: Comparing the DSRs of our Iteration-2 defense with the crawling control method given different LLMs.
@@ 5.4 Sensitivity Analysis
The placement of embedded policies within an HTML file had a significant effect on defense performance. Policies positioned at the top of a page yielded the highest DSR (up to 100%), compared to those placed mid-page (15%–25%) or at the bottom (5%–10%) (Figure 5). We infer that this pattern may be ascribed to the positional bias of LLMs, which tend to assign higher importance to tokens appearing earlier in LLM’s input sequence during generation Wang et al. (2025).
@@
Figure 6: Effect of policy visibility (visible as transparent webpage content vs. invisible as HTML meta tag) on DSRs across different LLMs.
@@ 7 Limitations
Despite the effectiveness of our proposed defensive framework, several limitations must be acknowledged. First, we primarily utilized fictitious webpages due to the practical constraints of conducting experiments on real-world websites, such as potential disruption to normal operations and limited availability of suitable real websites. This gap might cause discrepancies in evaluating how LLMs comply with embedded policies compared to real-world scenarios. Second, budget constraints limited our ability to perform extensive API interactions with advanced LLMs possessing sophisticated retrieval features. Our experiments were also constrained due to certain LLMs exhibiting unstable or unavailable web retrieval capabilities.
@@ A.2 Embedding Prompt
Embedding Prompt $z$ (Refusal to Answer & Visible & Baseline)
@@
Embedding Prompt $z$ (Refusal to Answer & Visible & Iteration 2)
@@
Embedding Prompt $z$ (Partial Mask & Invisible & Iteration 2)
@@
Embedding Prompt $z$ (Redirection & Invisible & Iteration 3)
@@ A.3 Synthetic Website Details
Table 5: List of synthetic websites used in experiments.
URLs with “*” denote redacted personal identifiers. URL Description Real homepages of individuals https://*1*.github.io/ A professional academic website presenting research interests, publications, teaching activities, and professional services. https://*2*.github.io/ A personal academic website showcasing research projects, publications, and demonstrations in various technical domains. Fictitious websites https://anonymous.4open.science/w/3-col-lab-EB2D/ A portfolio-style page featuring fictional tech projects in a three-column layout. https://anonymous.4open.science/w/education-1131/ An educational course showcase page featuring fictional design programs with highlights on art, UX, and visual communication. https://anonymous.4open.science/w/aerosky-BC89/ A mock real estate listing site featuring housing projects and property search tools. https://anonymous.4open.science/w/photograph-4591/ A photography portfolio site focused on portraits, wildlife, and visual storytelling. https://anonymous.4open.science/w/carcare-8F01/ A fictional EV company homepage featuring customizable vehicle services, smart integration, and battery innovations. https://anonymous.4open.science/w/creativeui-F6C7/ A tech company landing page offering fictional software, cloud, and app development services for digital transformation. https://anonymous.4open.science/w/portal-6DD9/ A mock news website presenting fictional headlines, featured articles, and blog content in a modern editorial layout. https://anonymous.4open.science/w/photoart-FC23/ A personal portfolio website for showcasing diverse photographic works. https://anonymous.4open.science/w/smartapp-2626/ A product landing page for a fake mobile app, with feature lists and app store badges. https://anonymous.4open.science/w/portfolio-6E5D/ A personal writing portfolio showcasing blog posts, copywriting skills, and storytelling projects.